10m of rope has a mass of 2.0kg. When the speed of transverse waves along the rope is 6.0ms^-1 the tension in N is?


(a) 1.2 (b) 7.2 (c) 72.0) (d) 28.8 (e) 2880 answers


1
Expert's answer
2022-02-28T13:52:25-0500

The speed of transverse waves along the rope is given by

"v=\sqrt{T/(m/l)}"

Hence, the tension in a string

"T=mv^2/l\\
=2.0*6.0^2/10=7.2\:\rm N"

Answer: (b)